JuniperBoi:

-17d - 6 + 18d= 2d - 8.... In this case, if I solved it correctly, d would equal 2, right?

justus:

Correct.

JuniperBoi:

ok just checking because I wasn't sure

justus:

No problem.

umm:

Given\[−17d−6+18d=2d−8\] Simplify both sides\[−17d−6+18d=2d−8\]\[−17d+−6+18d=2d+−8\]Combine like terms\[(−17d+18d)+(−6)=2d−8\]\[d+−6=2d−8\]\[d−6=2d−8\]Subtract 2d from both sides\[d−6−2d=2d−8−2d\]\[−d−6=−8\]Add 6 to both sides\[−d−6+6=−8+6\]\[−d=−2\]Divide both sides by -1\[\frac{ -d }{ -1 }=\frac{ -2 }{ -1 }\]Final answer\[d=2\]You are correct. (:

JuniperBoi:

thanks c:

umm:

No problem. Welcome to QuestionCove, by the way.

JuniperBoi:

Thanks. It's been really helpful, considering I'm a freshman just starting HS.

umm:

Great to hear, QuestionCove is a great learning tool for assistance. (:

JuniperBoi:

For sure, well have a nice day, or night c: